. Consider a function h(u,v)=4u^2+v^2. Find the level curves for k=n+1,n+2,n+3,n+4,n+5,n+6,where n is the sum of second and third digit of your arid number e.g.20-arid-470 take n=7+0=7. Perform all steps clearly. Also Sketch the neat Graph 


1
Expert's answer
2021-05-07T10:42:00-0400

Given, "h(u,v)=4u^2+v^2"


n=8 then

n+1


"\\4u^2+v^2=9,

\\4u^2+v^2=10,

\\4u^2+v^2=11,
\\
4u^2+v^2=12,\\

4u^2+v^2=13,
\\
4u^2+v^2=14,"


The graph is given as-